OpenStudy (anonymous):

Which inequality will have a solid boundary line and a shaded region above its graph? x − y ≥ 3 2x − 3y ≤ 3 3y − x < 2 2x + y < 7

OpenStudy (anonymous):

help??!!??!?!?!?

OpenStudy (texaschic101):

a solid line means there is an equal sign in the problem.That eliminates 2 of them right off. shaded above the line means it is greater then or to the right of the line. so we need to put the inequalities in y = mx + b form.. x - y >= 3 -y >= -x + 3 -- multiply by -1 to make x positive, and change the sign y <= x - 3.....this is less then...so the shading does not appear above the line. leaving us with : 2x - 3y <= 3 -3y <= -2x + 3 -- divide by -3 and change the sign y >= 2/3x - 1 ....this is greater then, so shading appears above the line.....so this is your answer.